OpenStudy (anonymous):

Solve the following equation. x- 3/4=2x+1/8 x=?

OpenStudy (anonymous):

Combine like Terms on one side then solve for x

OpenStudy (anonymous):

x-2x=1/4+3/4

OpenStudy (anonymous):

It looks like two "fractions" are set equal to each other. What method of solution is suggested?

OpenStudy (anonymous):

How'd they become 4s?

OpenStudy (anonymous):

That's your question : \[\color{Blue}{\Large x-\frac{3}{4}=2x+\frac{1}{8}}\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Do what i did in the first response

OpenStudy (anonymous):

When I solve it I get -7/8

OpenStudy (anonymous):

\[\Huge \color{Red}{That's~Correct~\bigstar}\]

OpenStudy (anonymous):

P=55003 Qd=530000-p/11 Qd=530000-55003/11 Qd=530000-5000 Qd=525000 is this correct?